OpenStudy (anonymous):

Given the formula, solve for the indicated variable. F=9/5c + 32 for c

OpenStudy (anonymous):

F = 9/5c + 32 ( to solve for c, u have to make c the subject) (take 32 to the left side) F-32 =9/5c (multiply all terms by 5) 5(F-32) = 9C divide all terms by 9 5/9 (F-32) = C

OpenStudy (anonymous):

how did 32 become a -32

OpenStudy (anonymous):

F=9/5c + 32 Subtract 32 from both sides F -32 =9/5c + 32 -32 F - 32 =9/5c

OpenStudy (anonymous):

@amber517 u subtract 35 from both side to make C alone F = 9/5c + 32 F - 32 = 9/5c + 32 -32 ( there u c 32 is getting cancelled in the right side, where u are left with C alone) so F - 32= 9/5c

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

DID U understand?

OpenStudy (anonymous):

yes